Zinc and cadmium have photoelectric work functions given by WZn=4.33eV and WCd=4.22eV, respectively. Calculate the maximum kinetic energy of photoelectrons from each surface if λ = 285 nm . Enter your answers numerically separated by a comma (KZn, KCd). Express your answer using two significant figures
Answer #1

Energy, E = 1243 eV/lambda = 1243/285 = 4.36 eV

Kinetic energy, KE = E - Wo

For Zinc, KE = 4.36 - 4.33 = 0.031 eV

For Cadmium, KE = 4.36 - 4.22 = 0.14 eV

(0.031, 0.14) eV

Comment in case any doubt please rate my answer ...